OpenStudy (anonymous):

Assume that, for a randomly chosen person,their next birthday is equally likely to occur on any day of the week,independently of any other person's birthday.Find the probability that,out of 350 randomly chosen people,at least 47 will have their next birthday on a monday.

OpenStudy (anonymous):

It's a binomial distribution.

OpenStudy (anonymous):

The probability of success (birthday on Monday) is \(\frac 17\).

OpenStudy (anonymous):

The total number of trials is \(350\).

OpenStudy (anonymous):

You could approximate it with a normal distribution. I feel that is what they want you to do here.

OpenStudy (anonymous):

5 mrks Anyoen could work it out for me :O I'll been with it for minutes -_-

OpenStudy (anonymous):

The mean would be \(\mu =np= 350\times \frac 17\) and the standard deviation would be \(\sigma =\sqrt{np(1-p)}=\sqrt{350\times \frac 17\times \frac 67}\).

OpenStudy (anonymous):

The \(z\) score would be \[ \frac {x-\mu}{\sigma }\]

OpenStudy (anonymous):

In this case \(x=47\)